# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 5.2.3 Identity
Identity, which pertains to the actors in a system, is another important aspect of context [122]. This includes the definition of individuals and their properties, such as name, address, gender, and age. For health monitoring, this can include additional information such as weight, height, and blood group. This is the most ubiquitous aspect of context in the systems, with nearly every system including personal information about the monitored individuals. Besides personal properties, identity also encompasses different user roles within the system. Most of the systems [76, 77, 78, 79, 80, 81, 83, 84, 85, 86, 89, 91, 92, 93, 94, 95, 98, 99, 100, 101, 104, 105, 106, 74, 107, 109, 70, 75, 111] support different users besides the individual being monitored, typically including clinicians and in some cases, caregivers and family members. Identity also includes agents, which are used in the agent-based systems [76, 80, 94, 95, 100, 108]. Agents [^27] An agent is a computer system situated in some environment that is capable of acting autonomously in order to achieve some goal(s) [125] have been applied extensively in the health domain [126] as well as in sensor-based systems [127]. The agent-based approach offers several advantages. For example, agents can be used as personal assistants to support humans in performing tasks and services [128]. This is explored in the system proposed by Mavropoulos et al. [100], which includes a smart virtual agent that clinicians can interact with via voice commands. Similarly, Yu et al. [70] implement an AI chatbot to answer user questions. Agent-based architectures and their advantages are discussed in greater detail in Section 7.
@@ 5.2.4 Activity
[…]
@@ 5.2.5 Other types of context
Besides location, time, identity, and activity, other types of contextual information are incorporated in the systems. Alti et al. [80] include hardware and network information as part of context, such as available communication protocols, CPU speed, battery power, and memory size. This information is used to ensure the efficient deployment of health services. Similarly, Zeshan et al. [111] use battery level and device response time to determine which caregiver or clinician’s device to send notifications to. Hristoskova et al. [92] incorporate media devices and their properties in their interpretation of context. For example, the screen size of devices such as mobile phones and tablets is used to determine how to display the health monitoring results. For small screens, the results are summarised. An important factor in health monitoring is the state of a person’s environment. A number of the systems use environmental data such as temperature and humidity from ambient sensors to provide additional context [76, 82, 85, 87, 96, 97, 98, 105, 107, 108, 71]. Weather data sources such as forecasts and indices are used by Kim et al. [97] to supplement sensor data, while the systems proposed by Akhtar et al. [76], Khozouie et al. [96], Vadillo et al. [108], and Zhou et al. [71] include sensors to monitor air quality by checking the levels of gases such as carbon monoxide, carbon dioxide, and oxygen. Contextual information can also include details about an individual’s diet, medication, and emotional state. These details are collected in the system proposed by De Brouwer et al. [83] through self-reporting via a mobile app. Similarly, Zafeiropoulos et al. [75] capture information about the user’s fatigue level.
@@ 5.2.6 The role of Semantic Web technologies
Ontologies appear to be particularly useful for the representation of contextual information, and majority of the selected systems reuse existing ontologies to do so. For instance, OWL-Time [^28] https://www.w3.org/TR/owl-time, an ontology that describes temporal properties of real-world objects such as sensors, is reused by a number of systems [83, 100, 105, 107, 110]. Friend of a Friend (FOAF) [^29] http://xmlns.com/foaf/0.1, an ontology that describes people profiles, is also widely reused among the systems [85, 90, 100, 73, 106, 107, 110]. Additionally, sensor ontologies, while not focused solely on contextual information, also include some aspects of context. For example, SAREF and SSN/SOSA ontologies include timestamps for sensor observations. Knowledge graphs can also be used to represent contextual information, as is done by Yu et al. [70], who use a knowledge graph to capture patient profile data from health records. Table 6 summarises the contextual information included in the systems and indicates which types of contextual information are captured using Semantic Web technologies.
@@
5 Alti et al. [80] L (patient, device); T (observation timestamps; report intervals); I (profile, user roles); O (hardware info, network info) L; T; I; O
@@
8 De Brouwer et al. [83] L (patient, headache location); T (observation timestamps, trigger duration); I (profile, user roles); A (sleep, physical activity, commute); O (diet, medication, mood) L; T; I; A; O
@@ 18 Hussain and Park [93] T (observation timestamps; time-domain features); I (profile, user roles) None
[…]
@@
L - location; T - time; I - identity; A - activity; O - other Table continued on next page.
Summary of contextual information captured in the systems and represented using Semantic Web technologies.
@@
41 Zeshan et al. [111] L (patient, caregiver, clinician); ; T (observation timestamps); I (profile, user roles); O (battery level; response time) L; T; I
@@
5.3 Situation analysis
A situation can be understood as a higher-level interpretation of sensor data that is relevant and of interest in an application domain [4]. Personal health monitoring systems should be capable of situation analysis. This entails both the detection and the prediction of health situations, which constitute two of the seven key challenges. We discuss each of them in turn.
5.3.1 Situation detection
In health monitoring systems, situation detection can take a variety of forms. One of these is the categorisation of individual sensor observations based on whether they are within or outside a given range as determined by domain knowledge. For example, in the system proposed by Akhtar et al. [76], when vital signs such as temperature and heart rate are outside the normal range, the situation is classified as an emergency. Likewise, Elhadj et al. [85] classify expected observations as normal, while observations outside the normal ranges are classified as abnormal. They also include a third classification, wrong, for faulty observations from malfunctioning sensors. Similar threshold-based situation categories are used in many of the systems [80, 83, 88, 89, 92, 94, 96, 99, 104, 105, 107, 109, 75, 111, 112]. Thresholds have also been used to classify physical activity based on level of intensity [81, 84, 86, 88, 94]. A better approach than using individual sensor observations is to consider different observations and personal attributes to classify individuals. This is done by Ali et al. [79], who classify the patient health condition as either healthy, moderate, or serious based on multiple sensor outputs and properties such as sex, weight, and height. Similarly, Chiang and Liang [82] classify situations as either healthy, moderate, or severe based on age, blood pressure, blood glucose, heart rate, and cholesterol.
[…]
With regard to techniques for situation detection, most of the systems implement some form of rule-based reasoning. Rules provide a way to implement expert knowledge in an if-then form, whereby if certain conditions are met, then a consequent conclusion is made or action taken. Despite their widespread use, rules have several limitations. Firstly, crisp rules are unable to handle uncertainty and ambiguity in sensor observations and the determination of health situations. To mitigate this, several systems incorporate fuzzy logic [79, 82, 86, 87, 103] and defeasible logic [76] in the rules. These techniques are discussed in greater detail in Section 5.6, which focuses on techniques for handling uncertainty in health monitoring. Secondly, rules are typically based on existing expert knowledge, and therefore cannot incorporate new knowledge that experts may be unaware of. Additionally, manually updating rules is time-consuming, making them difficult to scale. This challenge can be overcome using learned rules based on ML algorithms, which can acquire new, high quality knowledge automatically [130] and contribute to dynamic and adaptable rule-based systems. The systems proposed by Hussain et al. [93], Henaien et al. [90] and Peral et al. [104] extract rules from decision trees. However, caution should be exercised when using ML-derived rules, as they may still need verification and validation from domain experts. As an alternative to rule-based reasoning, Xu et al. [72] implement case-based reasoning, arguing that it is easier to capture human experiences using cases rather than rules. By searching for historical cases that are similar to the current case, their proposed system is able to obtain treatment plans that have been successful in the past.
In addition to the development of rules as discussed above, ML is also used in a number of the systems for the classification of diseases based not only on sensor data but also other data sources. Ali et al. [77] use a bidirectional long short-term memory (BiLSTM) model to detect diabetes and blood pressure, to classify sentiments from social networking data for mental health monitoring, and to classify drug side effects. Their proposed system uses domain ontologies to extract important features that can enhance the ML classification. Zhou et al. [71] also use a BiLSTM model for disease prediction. Other ML algorithms used include multi-layer perceptron for heart disease detection [78] and random forest for stroke detection [93]. ML is also used for physical activity classification, for example using the k-nearest neighbours [88, 100], decision trees [100], and random forest [94, 100] algorithms. Finally, ML can also be used to classify alert levels. This is done by Zafeiropoulos et al. [75], who use a graph neural network to distinguish between medium and high alerts. A full review of ML techniques for situation analysis in the health domain is outside the scope of this study. Readers are referred to the reviews by Ravì et al. [131] and Li et al. [132].
5.3.2 Situation prediction
A number of the selected systems explore the concept of risk as a situation prediction feature, since determining an individual’s risk profile for a certain condition can be used to predict future adverse health situations. For example, Alti et al. [80] use rules to determine the risk of death for diabetes patients based on high glucose levels and high heart rate. Rules are also used by Chiang and Liang [82] to determine the risk of arthritis recurring based on low temperatures and high humidity, and by Hristoskova et al. [92] also use rules to determine the risk of congestive heart failure (CHF) over a four-year time horizon. CHF risk stages are determined based on factors such as age, blood pressure, heart rate, and history of heart disease and diabetes. Similarly, in their use case of gestational diabetes, Rhayem et al. [105] use rules to determine the risk level for fetal loss based on age, presence of hyperglycemia, and presence of hypertension.
To support the identification of potential risks, future physiological readings can also be predicted using historic sensor observations, as is done by Peral et al. [104]. Their proposed system uses ML algorithms (support vector machine and logistic regression) to predict blood glucose levels over three-day and five-day windows. These predictions of sensor measurements can then be analysed to determine future health risks. Besides rules and ML, another technique used in some of the systems is Bayesian networks, which are probabilistic models in the form of directed acyclic graphs that can represent causal relationships among variables in a domain. Mchiek et al. [101] use a Bayesian network to calculate the risk of stroke occurring in the next seven days, based on risk factors such as age, presence of diabetes, high blood pressure, and symptom duration. This approach is also taken by Kordestani et al. [98] to determine the probability of the occurrence of kidney disease. The use of Bayesian networks and their contribution to explainability and uncertainty handling is discussed in greater detail in Sections 5.4 and  5.6 respectively.
As part of situation prediction, Zafeiropoulos et al. [75] explore the prediction of medication adherence using a graph neural network to identify potential lapses in pill dosing, while De Brouwer et al. [83] explore the detection of triggers as a means to anticipate potential headache attacks. Another useful aspect of situation prediction is the determination of the prognosis, i.e. expected progression, of a detected disease, although this is poorly explored in the systems. Hussain and Park [93] mention the intention to extend their system in future work to include automated stroke prognosis. In contrast, Yu et al. [110] include a disease progression class in their proposed ontology, representing past diagnoses or potential health risks and their associated times. However, the system does not include any methods to predict the progression of detected conditions.
@@ 5.3.3 The role of Semantic Web technologies
Semantic Web technologies can support situation analysis in two main ways: firstly, they formally represent important concepts and the relationships between them, i.e. sensor data, domain knowledge, contextual information, and even the situations themselves; and secondly, they support reasoning through which new knowledge can be derived from existing knowledge [4]. Although several situation-focused ontologies have been developed, including the Situation Theory Ontology [133] and the Scenes and Situations ontology [134], none of the selected systems extend any such ontologies. Despite this, Semantic Web technologies remain vital for situation analysis among the selected systems. Rule-based reasoning is the most common mechanism for situation analysis among the systems. These rules rely heavily on concepts that are formally defined in ontologies, and they are more often than not expressed in Semantic Web standard languages such as SWRL. The limitations of semantic-based approaches, such as scaling difficulties and inability to handle uncertainty, can be mitigated by combining them with complimentary techniques such as ML and Bayesian networks.
5.4 Decision Support
Decision support is the natural next step after situation analysis. Based on the detected and predicted situations, targeted support can be offered to mitigate adverse situations and promote favourable health outcomes.
5.4.1 Features and functionalities
Alerts and warnings are used in majority of the systems to warn of potentially dangerous situations and prompt mitigating action. These alerts are often sent to caregivers, clinicians, and emergency services depending on severity. However, several systems also take a patient-targeted approach, reminding the monitored individual about medications and exercise [82, 107, 108]. Alerts can also serve to remind users of medications, exercises, and medical tests. For example, the system proposed by Kordestani et al. [98] can remind clinicians to order additional laboratory tests when previously taken tests become out of date. A well-documented issue with alerts in the health domain is alert fatigue, a phenomenon in which users become desensitized to alerts due to their frequency [135]. Esposito et al. [86] mitigate this by differentiating between critical and non-critical abnormal situations, with the latter being sent out in a daily summary report email rather than an instantaneous notification for each case.
In addition to alerts, health monitoring systems can also trigger actions in response to adverse situations. For example, the system proposed by Alti et al. [80] triggers the injection of insulin in response to a blood glucose level above a certain threshold, while the system proposed by Hadjadj and Halimi [89] can trigger the opening of a vehicle door. Such systems must be integrated with an actuation device capable of carrying out the action. The system proposed by Titi et al. [107] includes several actuators such as a smoke alarm. Other systems are integrated with actuators capable of turning on lights and heaters [82], making emergency calls [105], or turning off water or gas if detected [108].
Another decision support feature is the generation of suggestions or recommendations for the mitigation of adverse situations. Several systems offer recommendations for lifestyle modifications, such as diet and exercise [78, 79, 81, 84, 97, 105, 106, 109, 70], as well as medication [84, 85, 92, 98, 104, 105, 107, 72]. An important factor when choosing appropriate treatment is the side effects of medications and how different medications interact with each other. Ali et al. [77] use drug review websites to collect data on side effects, while Elhadj et al. [85] keep track of medication interactions as well as patient allergies. This information assists clinicians in prescribing appropriate medications for each patient. Related to recommendations is the ability for the monitored individual to seek out relevant and trusted medical information. For example, the system proposed by Rhayem et al. [105] includes a notification module that allows patients to contact a clinician and receive recommendations and treatments from them.
5.4.2 Quality of decision support
Furthermore, the soundness of recommended interventions can be ensured by incorporating established and clinically validated medical guidelines. This can contribute to the acceptance of health monitoring systems by medical professionals and regulatory bodies. Despite this, few of the selected systems mention the use of such guidelines. They include Lopes de Souza et al. [99], who incorporate risk level classifications from the American Heart Association; De Brouwer et al. [83], who use the International Classification of Headache Disorders’ criteria to issue relevant alerts; and Ali et al. [78] and Hristoskova et al. [92], who the Framingham Risk Score to determine coronary heart disease risk.
5.4.3 The role of Semantic Web technologies
Similarly to situation analysis, the main value that Semantic Web technologies add to decision support in the selected systems is their ability to represent important concepts, particularly situations and domain knowledge, which can then be reasoned over. This is consistent with previous research findings on the use of Semantic Web technologies for decision support [138], with rules being the most common reasoning tool among the selected systems.
5.5 Explainability
A critical consideration in health monitoring systems is explainability, which contributes to the overall trustworthiness and adoption of such systems. For the purposes of this study, we adopt the perspective that explainability is essentially equivalent to interpretability [139, 140], which in turn can be defined as the degree to which a system’s operations can be understood by a human [141]. Two commonly used and complementary approaches to explainability in AI systems are: 1. prioritising human understanding of generated outputs; and 2. providing explicit explanations for those outputs [139].
5.5.1 Inherently interpretable techniques
One way to prioritise human understanding of system outputs is to use technologies and techniques that are inherently intuitive and easily comprehensible to humans. It can be argued that Semantic Web technologies meet this criteria and can contribute to the development of explainable systems; we discuss this in Section 5.5.3. However, the use of these technologies does not guarantee explainability. Other technologies and techniques implemented within the systems, and the ways in which they are combined, can also play a role in either enhancing or hindering the overall explainability of the system. For example, rules are inherently easy to understand [142], and nearly all the selected systems implement some form of rule-based reasoning. However, in some cases, rules are implemented for one aspect of the system while less interpretable techniques are used for other components. An instance of this is the system proposed by Ali et al. [78], where a deep learning model is used for disease prediction, while rule-based reasoning is applied for recommendation generation. This results in the decision support functionality being highly explainable, while the situation analysis component remains less so.
Bayesian networks can also be considered highly interpretable, as they can perform predictive and diagnostic reasoning [143] in a way that can be visually interpreted due to their graphical structure [144]. Predictive reasoning is done by Mcheick et al. [101], who use a Bayesian network to determine whether a person has a high risk of stroke based on risk factors. The reason for whether the risk is high or not can be traced back to the presence of risk factors. On the other hand, Kordestani et al. [98] use a Bayesian network for diagnostic reasoning, allowing for the understanding of a kidney disease diagnosis based on its causes. Fuzzy logic represents another class of interpretable techniques, since it allows variables and their classifications to be presented in a way that is intuitive [142]. Fuzzy approaches are used in a number of the selected systems; we discuss them in more detail in Section 5.6.1.
While ML is often criticized for its susceptibility to producing black box models, certain ML models are intrinsically interpretable, such as logistic or linear regression models and decision trees [145, 140]. However, even such models can be rendered uninterpretable as their complexity and scale increase, as is the case with large decision trees [145]. In cases where less interpretable ML models are used, they can be combined with Semantic Web technologies to enhance explainability. This is closely related to neuro-symbolic AI, in which the strengths of neural networks and symbolic methods are combined to achieve the best of both worlds [146]. A number of the selected systems explore such a hybrid approach, including Zafeiropoulos et al. [75] and Zhou et al. [71], who both use a knowledge graph to provide training data for deep learning models. Additionally, Ali et al. [77] use an ontology for feature extraction, providing some transparency into the selection of features for their BiLSTM model.
5.5.2 Post hoc explainability
The use of the use of inherently interpretable models should be prioritised in high-stakes domains [147]. Nonetheless, explaining the workings of a black box model after it has been trained is a well-established approach to explainability [140]. This can be done through methods such as the local interpretable model-agnostic explanations (LIME) technique [148], in which the black box model is approximated with an interpretable surrogate model that is then used to explain the original model’s predictions. However, such post hoc explainability methods are not mentioned in any of the selected systems.
Once the reasoning behind the system’s outputs has been determined, the next step is to communicate this to the users of the system. This explicit provision of explanations can also be considered a form of post hoc explainability. An explanation can be defined as the answer to a why-question [139], such as "Why was the situation classified as an emergency?", "Why was an alert sent to the clinician?", or "Why was this medication recommended?". A good explanation should have the ability to be contrastive (i.e. show the contrast between the situation chosen or decision made versus other possibilities), to distill a complex list of causes into one or two main ones, to highlight abnormalities, and to be general and probable [140]. The target audience is also an important consideration in the communication of explanations, since different audiences may value distinct aspects based on their role and perspective [149]. Health monitoring systems should therefore tailor explanations to the needs of various audiences, such as monitored individuals, clinicians, and caregivers. For certain audiences, like regulatory stakeholders, it may be more appropriate to include explanations as part of the overall system documentation. Only six of the selected systems report or indicate that some form of explanation is made available to users [76, 82, 85, 100, 57, 109, 70]. This is typically through a user interface, or, as in the system proposed by Akhtar et al. [76], as part of system logs. Chatbots and virtual agents, as implemented by Mavroppoulos et al. [100] and Yu et al. [70], can also provide explanations, since they are designed to answer queries from users.
5.5.3 The role of Semantic Web technologies
@@
Explainability relies on domain knowledge, and Semantic Web technologies excel at structuring such knowledge formally and unambigiously [150]. Ontologies can contribute to the development of explainable systems from three perspectives: by providing sound and explicit knowledge reference models; by supporting common-sense reasoning through the representation of context-aware semantic information; and by facilitating flexible knowledge abstraction and refinement [150]. Since most of the systems use ontologies to represent expert knowledge and contextual information, this can be seen as a first step towards achieving explainability. Additionally, many ontology reasoners provide explanations of the reasoning process, although it is not clear whether these explanations are made available to the end users. Knowledge graphs can also contribute to better understanding of system outputs in several ways, including providing a graph-based visualisation of concepts, entity and relation extraction from unstructured data, enrichment of datasets, and inference and reasoning [33]. While the connection between linked data and explainability is not direct, the open accessibility and interconnection of knowledge shared using a linked data approach can nonetheless contribute to explainability.

### v3

@@ 5.5.3 Identity
Identity, which pertains to the actors in a system, is another important aspect of context [155]. This includes the definition of individuals and their properties, such as name, address, gender, and age. For health monitoring, this can include additional information such as weight, height, and blood group. This is the most ubiquitous aspect of context in the systems, with every system including personal information about the monitored individuals. Besides personal properties, identity also encompasses different user roles within the system. Nearly all of the systems support different users besides the individual being monitored, typically including clinicians and in some cases, caregivers and family members, with the exception of 11 systems [88, 117, 92, 93, 121, 122, 125, 129, 101, 131, 133]. Identity also includes agents, which are used in the agent-based systems [109, 113, 108, 95, 96, 125, 126, 101]. Agents [^30] An agent is a computer system situated in some environment that is capable of acting autonomously in order to achieve some goal(s) [158] have been applied extensively in the health domain [159] as well as in sensor-based systems [160]. The agent-based approach offers several advantages. For example, agents can be used as personal assistants to support humans in performing tasks and services [161], such as the interactive virtual agent in the system proposed by Mavropoulos et al. [126]. Agent-based architectures and their advantages are discussed in greater detail in Section 7.
@@ 5.5.4 Activity
The fourth essential aspect of context is activity. This can refer to physical activity or the different activities of daily living such as eating and sleeping, both of which are important considerations for situation analysis. Activity can be derived from sensors such as accelerometers, or can be deduced from location or time (for example, a person in a bedroom in the middle of the night can be assumed to be sleeping). Physical activity is closely tied to health, and there are many physical activity guidelines issued by governments and global health organisations, including the World Health Organisation [162]. Due to this link between physical activity and health, 31 of the systems include physical activity as contextual information. Such systems monitor physical activity using smartphones, smart watches, or inertial measurement units, which combine accelerometers, gyroscopes, and in some cases, magnetometers [91] [110, 111, 115, 116, 90, 92, 95, 96, 121, 124, 126, 129, 107]. Chiang and Liang [88] monitor body movement using motion sensors placed around the home. This serves two purposes. Firstly, the individual’s movement within the home is able to be monitored. This can determine their location at any given time. Secondly, they are able to interact with the system using body movements, such as hand-waving to activate the system. Ali et al. [112] similarly use motion sensors to keep track of body movement. They use range of motion as a metric, which is particularly important for elderly patients who may lose their ability to perform daily activities as their range of motion decreases. The systems proposed by Ivașcu et al. [96] and Zafeiropoulos et al. [107] perform gait analysis, capturing features such as freezing of gait, postural instability, and rigidity.
Self-reported information can also be used to determine physical activity, but this may not be accurate. To mitigate this, Chatterjee et al. [114] use a combination of sensor and questionnaire data. Sensors are used to monitor number of steps and duration of activity, while questionnaires are used to determine the type of activity, for example running or weightlifting. Beyond tracking physical activity, activity recognition is also important in health monitoring. It can help in the detection of adverse events like falls, as is done in the systems proposed by Chiang and Liang [88], Vadillo et al. [101], and Zafeiropoulos et al. [107]. Additionally, seven systems are able to recognise daily activities such as sitting, walking, and sleeping [115, 91, 92, 95, 126, 83, 107].
@@ 5.5.5 Other types of context
Besides location, time, identity, and activity, other types of contextual information are incorporated in the systems. Alti et al. [113] include hardware and network information as part of context, such as available communication protocols, CPU speed, battery power, and memory size. This information is used to ensure the efficient deployment of health services. Similarly, Zeshan et al. [132] use battery level and device response time to determine which caregiver or clinician’s device to send notifications to. Hristoskova et al. [119] incorporate media devices and their properties in their interpretation of context. For example, the screen size of devices such as mobile phones and tablets is used to determine how to display the health monitoring results. For small screens, the results are summarised. An important factor in health monitoring is the state of a person’s environment. 15 of the systems use weather data such as temperature and humidity from ambient sensors to provide additional context [109, 87, 114, 88, 89, 117, 91, 93, 121, 122, 123, 125, 83, 100, 101, 103]. Weather data sources such as forecasts and indices are used by Kim et al. [122] to supplement sensor data, while the other systems [109, 121, 125, 101, 103] include sensors to monitor air quality by checking the levels of different gases and/or inhalable particulate matter in the air. Contextual information can also include details about an individual’s diet, medication, and emotional state. These details are collected in the system proposed by De Brouwer et al. [115] through self-reporting via a mobile app. Finally, Martella et al. [125] and Zafeiropoulos et al. [107] capture information about the user’s fatigue level.
@@ 5.5.6 The role of Semantic Web technologies
Ontologies appear to be particularly useful for the representation of contextual information, and majority of the selected systems reuse existing ontologies to do so. For instance, OWL-Time [^31] https://www.w3.org/TR/owl-time, an ontology that describes temporal properties of real-world objects such as sensors, is reused by a number of systems [115, 126, 83, 100, 131]. Friend of a Friend (FOAF) [^32] http://xmlns.com/foaf/0.1, an ontology that describes people profiles, is also widely reused among the systems [108, 91] [89, 93, 126, 106, 98, 100, 131]. Additionally, sensor ontologies, while not focused solely on contextual information, also include some aspects of context. For example, SAREF and SSN/SOSA ontologies include timestamps for sensor observations. Knowledge graphs can also be used to represent contextual information, as is done by Ammar et al. [108] and Yu et al. [104], who use a knowledge graph to capture patient profile data from health records. Linked data is also useful for context awareness. For instance, Ammar et al. [108] use linked open data to access global knowledge from the web, including data relating to social determinants of health. Additionally, Martella et al. [125] mention the possibility of using linked open data as a source of external information, although this is not implemented in the system.
Table LABEL:context-table summarises the contextual information included in the systems and indicates which types of contextual information are captured using Semantic Web technologies.
@@
5 Alti et al. [113] L (patient, device); T (observation timestamps, report intervals); I (profile, user roles); O (hardware info, network info) L; T; I; O
6 Ammar et al. [108] L (patient); I (profile, user roles); A (exercise); O (social determinants of health) L; I; A; O
7 Bampi et al. [87] L (patient); T (observation timestamps & latency); I (profile, user roles); O (weather) L; T; I; O
@@
10 De Brouwer et al. [115] L (patient, headache location); T (observation timestamps, trigger duration); I (profile, user roles); A (sleep, physical activity, commute); O (diet, medication, emotion) L; T; I; A; O
@@
15 Garcia-Moreno et al. [91] L (activity); T (observation timestamps, duration, sampling frequency, time-domain features); I (profile, user roles); A (recognition) L; T; I; A
@@ 21 Hussain and Park [120] T (observation timestamps; time-domain features); I (profile, user roles) None
22 Ivașcu and Negru [95] T (time-domain features); I (profile, user roles); A (recognition, intensity level) T; I; A
23 Ivașcu et al. [96] T (observation timestamps); I (profile, user roles); A (sleep quality, gait analysis) A
@@
25 Kilintzis et al. [97] T (observation timestamps); I (profile, user roles); A (step count, sleep stage ratio) T; I; A
@@
29 Martella et al. [125] L (patient); T (observation timestamps & duration); I (profile); A (physical effort); O (air quality, weather, ambient light, sound pressure, electromagnetic emissions, fatigue) L; T; I; A; O
@@
46 Zeshan et al. [132] L (patient, caregiver, clinician); T (observation timestamps); I (profile, user roles); O (battery level; response time) L; T; I
@@
5.6 Explainability
A critical consideration in health monitoring systems, particularly those that incorporate AI, is explainability. For the purposes of this study, we adopt the perspective that explainability is essentially equivalent to interpretability [163], which in turn can be defined as the degree to which a system’s operations can be understood by a human [164]. Explainability contributes significantly to the overall trustworthiness and adoption of systems in the health domain. Two complementary approaches to explainability in AI systems are prioritising human understanding of generated outputs, and providing explicit explanations for those outputs [163]. The former can be achieved by using techniques that are inherently intuitive and easily comprehensible (intrinsic explainability), or by explaining the workings of a model after it has been trained (post hoc explainability) [165]. We examine both intrinsic and post hoc explainability below, followed by a consideration of explicit explanations and their quality.
5.6.1 Intrinsic explainability
It can be argued that Semantic Web technologies are inherently intuitive and can contribute to the development of explainable systems; we discuss this in Section 5.6.4. However, the use of these technologies does not guarantee the explainability of the system as a whole. Other technologies and techniques implemented within the systems, and the ways in which they are combined, can also play a role in either enhancing or hindering the overall explainability of the system. For example, rules are inherently easy to understand [166], and nearly all the selected systems implement some form of rule-based reasoning using Semantic Web rule languages. However, in some cases, rules are implemented for one aspect of the system while less interpretable techniques are used for other components. An instance of this is the system proposed by Ali et al. [111], where a deep learning model is used for disease prediction, while rule-based reasoning is applied for recommendation generation. This results in the decision support functionality being highly explainable, while the situation prediction component remains less so.
Bayesian networks can also be considered highly interpretable, as they can perform predictive and diagnostic reasoning [167] in a way that can be visually interpreted due to their graphical structure [168]. Predictive reasoning is done by Mcheick et al. [127], who use a Bayesian network to determine whether a person has a high risk of stroke based on risk factors. The reason for whether the risk is high or not can be traced back to the presence of risk factors. On the other hand, Kordestani et al. [123] use a Bayesian network for diagnostic reasoning, allowing for the understanding of a kidney disease diagnosis based on its causes. Fuzzy logic represents another class of interpretable techniques, since it allows variables and their classifications to be presented in a way that is intuitive [166]. Fuzzy approaches are used in five of the selected systems; we discuss them in more detail in Section 5.7.1.
While ML is often criticized for its tendency to produce black box models, certain ML models are intrinsically interpretable, such as logistic or linear regression models and decision trees [169, 165]. However, even such models can be rendered uninterpretable as their complexity and scale increase, as is the case with large decision trees [169]. In cases where less interpretable ML models are used, they can be combined with Semantic Web technologies to enhance explainability. This is closely related to neuro-symbolic AI, in which the strengths of neural networks and symbolic AI are combined to achieve the best of both worlds [170]. Zafeiropoulos et al. [107] and Zhou et al. [103] explore this hybrid approach, with both using a knowledge graph to provide training data for deep learning models. Additionally, Ali et al. [110] use an ontology for feature extraction, providing some transparency into the selection of features for their BiLSTM model.
5.6.2 Post hoc explainability
The use of inherently interpretable models should be prioritised in high-stakes domains [171]. Nonetheless, explaining the workings of a model after it has been trained is a well-established approach to explainability [165]. Two well-known post hoc explainability techniques are Local Interpretable Model-agnostic Explanations (LIME) [172] and SHapley Additive exPlanations (SHAP) [173]. LIME approximates a black box model with an interpretable surrogate model that is then used to explain the original model’s predictions. SHAP, by contrast, uses Shapley values from coalitional game theory to explain individual predictions by quantifying how much each feature contributes to the outcome. Beyond these methods, plotting techniques can also be used to visualise feature effects. They include partial dependence plots, which show the marginal effects of one or two features on predictions; accumulated local effects plots, which provide more accurate feature effects by accounting for feature correlations; and individual conditional expectation plots, which show how an individual prediction changes when a feature changes [165]. However, such post hoc explainability methods are not mentioned in any of the selected systems.
5.6.3 Explanation quality
Once the reasoning behind the system’s outputs has been determined, the next step is to communicate this to the users of the system through explicit explanations. An explanation can be defined as the answer to a why-question [163], such as "Why was the situation classified as an emergency?", "Why was an alert sent to the clinician?", or "Why was this medication recommended?". Good explanations have the following characteristics [165]:
  1. 1.
    They are contrastive, showing the contrast between the selected outcome versus other possibilities.
  2. 2.
    They are selected, distilling long lists of causes into one to three main ones.
  3. 3.
    They highlight abnormalities, specifying causes that had a small probability but nevertheless occurred and influenced the outcome.
  4. 4.
    They are consistent with existing knowledge or prior beliefs, as humans tend to devalue or ignore explanations that differ from their beliefs due to confirmation bias. This is especially important for clinicians, since explanations that contradict established medical knowledge can undermine their trust.
  5. 5.
    They are general and probable, identifying causes that can explain many similar outcomes.
The target audience is also an important consideration in the communication of explanations, since different audiences may value distinct aspects based on their role and perspective [174]. Health monitoring systems should therefore tailor explanations to the needs of various audiences, such as monitored individuals, clinicians, and caregivers. For certain audiences, like regulatory stakeholders, it may be more appropriate to include explanations as part of the overall system documentation. Only eight of the selected systems report or indicate that some form of explanation is made available to users [114, 125] [109, 88, 89, 126, 83, 102, 104]. This is typically through a user interface, messaging platform, or, as in the system proposed by Akhtar et al. [109], as part of system logs. Chatbots and virtual agents, as implemented by Mavroppoulos et al. [126] and Yu et al. [104], can also provide explanations, since they are designed to answer queries from users.
@@ 5.6.4 The role of Semantic Web technologies
Domain knowledge can enhance explainability [175], and Semantic Web technologies excel at structuring such knowledge formally and unambiguously [176]. Ontologies can contribute to the development of explainable systems from three perspectives: by providing sound and explicit knowledge reference models; by supporting common-sense reasoning through the representation of context-aware semantic information; and by facilitating flexible knowledge abstraction and refinement [176]. Since most of the systems use ontologies to represent expert knowledge and contextual information, this can be seen as a first step towards achieving explainability. Moreover, many ontology reasoners such as HermiT [177] and Pellet [178] provide justifications for their inferences. These justifications can be generated into human-readable explanations, which in turn can be presented to system users [179]. Although 11 systems [112, 114, 116, 89, 90, 94, 119, 121, 100, 101, 107] mention using ontology reasoners, none of them indicate that the reasoner justifications are presented to users as explanations. Beyond reasoners, explanations themselves can be represented using ontologies. For instance, the Explanation Ontology [180] is a general-purpose ontology that connects explanations to underlying data and knowledge. Similarly, the Evidence and Conclusion Ontology [181] captures evidence to support annotations and assertions in the biomedical domain, which can be used to generate explanations. Yet, none of the reviewed systems formalize explanations in this way.
Knowledge graphs can also contribute to better understanding of system outputs in several ways, including providing a graph-based visualisation of concepts, entity and relation extraction from unstructured data, enrichment of datasets, and inference and reasoning [46]. Finally, although the connection between linked data and explainability is not direct, the open accessibility and interconnection of knowledge shared using a linked data approach can nonetheless contribute to explainability.